OpenStudy (anonymous):

Solve the following system using the substitution method. 3x - y = -2 6x + 5y = 24

OpenStudy (anonymous):

Cool, so what you want to do is, isolate one of the variables ( x or y ) meaning get it on it's own, and then substitute it into the second equation.

OpenStudy (anonymous):

So lets work with the first expression, do you want to get y on its own or x?

OpenStudy (anonymous):

I don't care about the answers, I'm trying to help you figure it out, so do you want to learn this?

OpenStudy (anonymous):

Or are you just looking for the answers.

OpenStudy (anonymous):

i am not good english so it hard to explan

OpenStudy (anonymous):

Ok but I'm asking you, are you looking for just the answer or the explanation, so a) answer only b) explanation Pick either a) or b)

OpenStudy (anonymous):

b

OpenStudy (anonymous):

Alright, good, so what would you do given the information above?

OpenStudy (anonymous):

solve using substitution methud?

OpenStudy (anonymous):

Ok, I'll ask you this, can you find x for me here: \[3x-y=-2\] Solve for x.

OpenStudy (anonymous):

3x=y-2 ?

OpenStudy (anonymous):

Yes, you're on the right track, but you want x on its own, so can you finish that problem off? Meaning can you get 3 on the right side.

OpenStudy (anonymous):

x=y-5?, sorry this is a bit hard for me

OpenStudy (anonymous):

Um no you have to divide by 3, remember what you do to one side you must do to the other, and it's fine if it's hard for you right now. Don't worry about making mistakes, that's how we learn, and also I don't judge. SO here is how you would solve for x.  Does that make sense?